(1) Let Q(x, y) be the statement "x+y=xy." If the domain for both variables consists of all integers, what are the truth values?
(a)Q(1,1)
(b)Q(2,0)
(c)∀yQ(1, y)
(d)∃xQ(x,2)
(e)∃x∃yQ(x, y)
(f)∀x∃yQ(x, y)
(g)∃y∀xQ(x, y)
(h)∀y∃xQ(x, y)
(i)∀x∀yQ(x, y)
1
Expert's answer
2020-09-09T19:35:53-0400

a. Q(1,1)


"x+y=1+1=2""xy=(1) (1)=1"

Since x+y is not equal to xy, truth value of Q(x, y) is FALSE

Answer: FALSE


b. Q(2,0)


"x+y=2+0=2""xy=2*0=0"

Since x+y is not equal to xy, the truth value of Q(x, y) is FALSE

Answer: FALSE


c. "\forall yQ(1,y)"

By contradiction;

let y be 1:


"x+y=1+1=2""xy=1*1=1"

Since x+y is not equal to xy, the truth value of Q(x, y) is FALSE

Answer: FALSE


d. "\exist xQ(x, 2)"

let x be 2:


"x+y=2+2=4""xy=2*2=4"

Since x+y=xy, the truth value of Q(x, y) is TRUE

Answer : TRUE


e. "\exist x\exists yQ(x, y)"

let x=2 and y=2


"x+y=2+2=4""xy=2*2=4"

Since x+y=xy, the truth value of Q(x, y) is TRUE

Answer : TRUE


f. "\forall x \exists yQ(x, y)"

Let x=1 and y=1:


"x+y=1+1=2""xy=1*1=1"

Since x+y is not equal to xy, then the truth value of Q(x, y) is FALSE

Answer : FALSE


g. "\exist y \forall xQ(x, y)"

Let x=1 and y=1:


"x+y=1+1=2""xy=1* 1=1"


Since x+y is not equal to xy, then the truth value of Q(x, y) is FALSE


Answer : FALSE


h. "\forall y \exist xQ(x, y)"

Let x=1, y=1:


"x+y=1+1=2""xy=1* 1=1"

Since x+y is not equal to xy, then the truth value of Q(x, y) is FALSE


Answer : FALSE


i. "\forall x\forall yQ(x, y)"

Let x=1, y=1:


"x+y=1+1=2""xy=1*1=1"

Since x+y is not equal to xy, then the truth value of Q(x, y) is FALSE


Answer : FALSE